OpenStudy (anonymous):

Given the following system of equations, identify the type of system. y = -x + 4 2x + 2y = 8 consistent inconsistent equivalent

pooja195 (pooja195):



pooja195 (pooja195):

Can you tell what the answer is?

OpenStudy (anonymous):

inconsistent? @pooja195

OpenStudy (math&ing001):

Lets put them both in standard form: y = -x + 4 <=> x+y=4 (2x + 2y = 8)/2 <=> x+y=4 You can see that they actually represent the same line (aka equivalent).

OpenStudy (anonymous):

thank you:)